# EducationQ: Evaluating LLMs’ Teaching Capabilities Through Multi-Agent Dialogue Framework

### v1

@@ Abstract
Large language models (LLMs) increasingly serve as educational tools, yet evaluating their teaching capabilities remains challenging due to the resource-intensive, context-dependent, and methodologically complex nature of teacher-student interactions. We introduce EducationQ, a multi-agent dialogue framework that efficiently assesses teaching capabilities through simulated dynamic educational scenarios, featuring specialized agents for teaching, learning, and evaluation. Testing 14 LLMs across major AI Organizations (OpenAI, Meta, Google, Anthropic, and others) on 1,498 questions spanning 13 disciplines and 10 difficulty levels reveals that teaching effectiveness does not correlate linearly with model scale or general reasoning capabilities - with some smaller open-source models outperforming larger commercial counterparts in teaching contexts. This finding highlights a critical gap in current evaluations that prioritize knowledge recall over interactive pedagogy. Our mixed-methods evaluation, combining quantitative metrics with qualitative analysis and expert case studies, identifies distinct pedagogical strengths employed by top-performing models (e.g., sophisticated questioning strategies, adaptive feedback mechanisms). Human expert evaluations show 78% agreement with our automated qualitative analysis of effective teaching behaviors, validating our methodology. EducationQ demonstrates that LLMs-as-teachers require specialized optimization beyond simple scaling, suggesting next-generation educational AI prioritize targeted enhancement of specific pedagogical effectiveness [^1] Y. Shi, R. Liang contributed equally to this work. [^2] https://github.com/SunriserFuture/EducationQ.
@@
Yao Shi1,2, Rongkeng Liang2, Yong Xu1, 1South China University of Technology, 2Education Innovation Research Institute of Guangdong, {yao, nico}@sunriser.org, yxu@scut.edu.cn
@@ 1 Introduction
Existing LLM evaluation approaches - whether through closed-ended questions, open-ended responses, or multi-turn dialogues - present fundamental limitations in assessing teaching capabilities. Current benchmarks predominantly rely on closed-ended assessments, which enable efficient automation but fail to capture the complexity and teacher agency in educational interactions. While open-ended evaluation could better reflect teaching dynamics, it faces significant challenges in scalability and consistency due to reliance on human judgment. Multi-turn dialogue frameworks, despite better capturing interactive complexity, lack specific mechanisms for evaluating teaching effectiveness. These limitations particularly impact teaching evaluation: benchmarks neither capture teachers’ active role in questioning, assessment, and real-time adaptation, nor provide scalable solutions for assessing teaching quality.
To address these challenges, we propose EducationQ, a novel multi-agent dialogue framework that incorporates formative assessment into the evaluation of LLMs’ teaching capabilities. Formative assessment—a continuous process of evaluating learner progress, identifying gaps, and adjusting teaching strategies William (2011)—is essential for personalized instruction. It bridges the gap between current abilities and potential, enhances learning outcomes, and promotes educational equity through AI Pardo et al. (2019); Ruiz-Primo and Furtak (2007); U.S. Department of Education, Office of Educational Technology (2023); Allal and Pelgrims Ducrey (2000). In classroom settings, informal formative assessments (IFAs) are a common practice during instructional dialogues, where teachers pose questions, assess student understanding, and provide timely feedback and guidance Sezen-Barrie and Kelly (2017); Guskey (2005).
The EducationQ framework models these interactions through a triadic system of teacher, student, and evaluator agents, simulating cyclical teacher-student interaction. This design captures teachers’ agency in employing diverse strategies and navigating complex educational contexts while enabling automated evaluation of dialogue quality. To support this framework, we curated a robust dataset of 1,498 questions from established benchmarks GPQA and MMLU-Pro, spanning diverse disciplines and difficulty levels. We employed a mixed-methods approach to comprehensively evaluate LLMs’ teaching capabilities. Our approach quantifies teaching effectiveness from an outcome-aligned perspective Gitomer and Duschl (2007) by measuring student learning gains, while analyzing pedagogical strategies using an automated qualitative evaluator agent with strong alignment (78% agreement) with human education experts.
@@
  • A multi-agent dialogue methodology for simulating and assessing authentic teaching interactions.
@@ 2.1 LLM Evaluation
Task-oriented performance benchmarks like MMLU Hendrycks et al. (2021a), MMLU-Pro Wang et al. (2024b), and GPQA Rein et al. (2023) employ closed-ended questions to evaluate domain knowledge and reasoning abilities. Similarly, MATH Hendrycks et al. (2021b) examines mathematical reasoning, while HumanEval Chen et al. (2021) tests programming capabilities. Recent work by Liang et al. (2024) introduces MathChat, which evaluates LLMs on multi-turn mathematical interactions, revealing that models often struggle with sustained reasoning across multiple dialogue turns despite strong single-turn performance.
@@ 2.3 LLM-Based Student Modeling
Recent work has explored using LLMs to simulate student behavior and interactions. Xu & Zhang (2023) investigated the feasibility of using generative students to test educational materials. Markel et al. (2023) employed LLMs to simulate student dialogues for teacher training. Lu & Wang (2024) found that profile-based generative students closely mirror human student performance in MCQ responses. Jin et al. (2025) proposed TeachTune, a framework generating pedagogical agent dialogues with diverse simulated student profiles for human evaluation, complementing our automated fixed-student-model assessment approach.
@@ 3 Dataset
We constructed our evaluation datasets, as summarized in Table 1, by systematically curating questions from two well-established benchmarks: GPQA (n=448), featuring domain expert-authored questions, and MMLU-Pro (n=12,032), containing reasoning-intensive questions with enhanced robustness through 10-option design across 14 educational categories. These datasets span undergraduate to PhD-level content, providing a rigorous foundation for evaluating teaching capabilities. The resulting dataset benefits from the well-established credibility of source benchmarks, proven stability of their evaluation protocols, and comprehensive representation across multiple teaching dimensions.
@@
Figure 2: Dataset distribution across 13 academic disciplines and 10 difficulty levels. Left: original MMLU-Pro shows uneven distribution across subjects and difficulty levels; Right: MMLU-Pro Stratified with our balanced sampling approach ensuring equal representation.
@@ 4.2 Teacher Agent
To prevent direct answer disclosure, we implement strict data flow controls that technically prevent teacher agents from accessing question options. This enforced constraint, combined with explicit prompt instructions, ensures teachers must guide learning without revealing answers.
@@ 4.3 Evaluator Agent
The evaluator agent is initiated and (see Appendix A.3) as an education assessment expert well-versed in pedagogical theory and practice. It validates the dialogues and evaluates teaching dialogues according to specified dimensions and compares teacher performances to determine superior approaches.
The assessment framework comprises 17 distinct scoring dimensions, including teacher-focused metrics (questioning, assessment, feedback) and student-impact measures (metacognitive reflection, knowledge dimension, etc) Krathwohl (2002); Looney (2011); Wilen (1987); Wass and and (2014). Given the exploratory nature of this component, we did not address dimensional overlap.
@@ 4.5 Pre-Test
The pre-test establishes the student agent’s initial knowledge baseline while providing teachers with preliminary insights through chain-of-thought reasoning patterns. To ensure broader applicability and stability, we conducted pre-test evaluation following official MMLU-Pro and GPQA Diamond benchmark protocols and parameters.
@@ 4.8 Content Boundary Design
As previously mentioned in the Teacher Agent section, to prevent direct answer disclosure, we implemented the following technical constraints: (a) Teacher agents cannot access any of the answer options, relying solely on student reasoning patterns and correctness judgments for guidance. (b) Students cannot access pre-test correctness judgments during dialogue, learning exclusively through teacher interaction. (c) Students retain access to complete question content including options, enabling learning through experience association.
@@ 4.9 Interaction Parameters and Constraints
Ablation studies identified optimal parameters of 150 tokens per dialogue round across five rounds, balancing effectiveness and efficiency. Increasing token limits to 250 yielded no significant learning gains, while reducing teacher dialogue to 70-100 tokens degraded teaching performance. Doubling rounds to 10 (with halved student token limits) increased computational costs without surpassing the effectiveness of the 5-round, 150-token configuration. In final experiments, teacher responses averaged 73.6 tokens, with student responses averaging 260 tokens.
@@ 7.2 Framework Robustness and Stability
Table 3: Stability Study of the Multi-agent Framework
@@
Table 4: Teaching Performance Comparison of Large Language Models
@@ 7.5 Teaching Stability Analysis
Through analysis of the Positive-Negative Impact Ratio (PNIR, Equation 2, lower is better), we identified significant variations in teaching stability across models. Llama 3.1 70B Instruct demonstrated exceptional stability, generating only 36 negative cases against 200 positive improvements (PNIR = 0.18). While Gemini 1.5 Pro 002 achieved comparable positive cases (188), its higher PNIR of 75 indicated greater performance volatility. OpenAI o1-mini and Llama 3.1 405B Instruct maintained moderate stability (PNIR = 0.25 and 0.28 respectively). These findings suggest that high teaching effectiveness and high stability can coexist.
@@ 8.1 Teaching Strategy Analysis: Expert-Annotated Case Studies
Human education experts who reviewed these dialogues confirmed that learning gains resulted from diverse and sophisticated teaching strategies rather than direct answer disclosure. The experts identified and annotated clear evidence of pedagogical techniques. Our expert-annotated case studies provide detailed evidence of how different LLMs implement distinct teaching strategies and demonstrate varying pedagogical skills and explain why models with superior general capabilities might underperform in educational interactions.
@@ 9 Conclusion
These findings suggest a fundamental rethinking of educational LLM development: prioritizing specialized teaching capabilities over general model scaling. The significant performance variations also highlight the inadequacy of traditional LLM metrics for predicting teaching effectiveness, underscoring the critical need for specialized, interaction-based evaluation frameworks like EducationQ to guide the development of effective AI in education.
@@
Our decision not to extensively use evaluators’ evaluations or incorporate qualitative indicators in primary benchmark testing meant our main metrics might not fully capture the diversity of LLMs’ teaching capabilities. While evaluator-based analysis provided valuable insights, its key findings require further validation.
@@ A.1 Student Agent Configuration
The student agent uses a consistent prompting template designed to simulate authentic student learning behavior, including expressing uncertainty and focusing on problem analysis. Below we detail the exact configuration and prompting templates used in our experiments for the designated student agent.
@@ A.1.3 Response Format
For dialogue interactions, The student agent’s interaction involves receiving context from previous rounds and generating a response for the current round:
@@
Note: [SYSTEM_MESSAGE] refers to the template shown in Appendix A.1.2. [PRE_TEST_INFO] is included in the context if the corresponding configuration parameter (include_pretest_info) is enabled.
@@ A.2.3 Pre-test Information Format
Note: The [CORRECTNESS] is either "Correct." or "Incorrect." based on the student’s pre-test selection.
@@ A.3.3 Evaluation Dimensions
The evaluator assesses teaching effectiveness based on dimensions specific to each analytical perspective:
@@ A.3.4 Evaluation Input Format (Example: Holistic Interaction Analysis)
Note: [FORMATTED_INTERACTION] includes Teacher: and Student: turns. For Teacher/Student-centric analysis, only the relevant parts (questions or answers) are included.
@@ A.3.5 Evaluation Output Format (Each Dimension)
For each dimension within a specific analysis task, the evaluator provides a JSON object structured as follows (refer to Appendix B.3 for scoring guidelines):
@@
(This structure is nested within the schemas shown in the next subsection.)
@@ A.3.6 Comparative Analysis Output Format
Note: The code includes logic to de-anonymize the teacher_a/teacher_b keys and the final choice based on the actual teacher models being compared.
The implementation details, including the specific JSON schemas passed to the API for each analysis type, can be found in the EvaluatorLLM class.
@@ B.2 Quality Control and Response Validation
Several mechanisms ensure the quality and validity of LLM responses during generation and processing:
@@ B.3 Evaluator Scoring Guidelines
All qualitative scoring performed by the Evaluator agent (Appendix A.3.2) follows a standardized 1-10 scale:
@@ Appendix C Model Specifications
This section provides detailed specifications for the language models evaluated in this study. Table 7 lists the models, their developers (Organization), the platform or API provider through which they were accessed, their data precision type, context window size, and approximate parameter count.
@@
  • Note: "-" indicates unspecified information. Context: Context window sizes are in tokens. Org.: Organization (model developer). Provider: serving platform.
@@ Appendix D Dataset Distribution
Table 8 presents the complete distribution of our evaluation datasets across different disciplines and sources.
@@ 14 GPQA
Table 8: Distribution of Evaluation Datasets by Source and Discipline
@@ Appendix E Teaching Performance Ranking by Subject
Tables 9 through 21 present the detailed Absolute Learning Gain ($\Delta$, in %) for each teacher model within 13 academic disciplines. Models are ranked based on their $\Delta$.

### v2

@@ Abstract
Large Language Models (LLMs) increasingly serve as educational tools, yet evaluating their teaching capabilities remains challenging due to the resource-intensive, context-dependent, and methodologically complex nature of teacher-student interactions. We introduce EducationQ, a multi-agent dialogue framework that efficiently assesses teaching capabilities through simulated dynamic educational scenarios, featuring specialized agents for teaching, learning, and evaluation. Testing 14 LLMs across major AI Organizations (OpenAI, Meta, Google, Anthropic, and others) on 1,498 questions spanning 13 disciplines and 10 difficulty levels reveals that teaching effectiveness does not correlate linearly with model scale or general reasoning capabilities - with some smaller open-source models outperforming larger commercial counterparts in teaching contexts. This finding highlights a critical gap in current evaluations that prioritize knowledge recall over interactive pedagogy. Our mixed-methods evaluation, combining quantitative metrics with qualitative analysis and expert case studies, identifies distinct pedagogical strengths employed by top-performing models (e.g., sophisticated questioning strategies, adaptive feedback mechanisms). Human expert evaluations show 78% agreement with our automated qualitative analysis of effective teaching behaviors, validating our methodology. EducationQ demonstrates that LLMs-as-Teachers require specialized optimization beyond simple scaling, suggesting next-generation educational AI prioritize targeted enhancement of specific pedagogical effectiveness [^1] Y. Shi, R. Liang contributed equally to this work., [^2] https://github.com/SunriserFuture/EducationQ.
@@
Yao Shi1,2, Rongkeng Liang2, Yong Xu1 1South China University of Technology, 2Education Innovation Research Institute of Guangdong {yao, nico}@sunriser.org, yxu@scut.edu.cn
@@ 1 Introduction
Existing LLM evaluation approaches - whether through closed-ended questions, open-ended responses, or multi-turn dialogues - present fundamental limitations in assessing teaching capabilities. Current benchmarks predominantly rely on closed-ended assessments, which enable efficient automation but fail to capture the complexity and teacher agency in educational interactions. While open-ended evaluation could better reflect teaching dynamics, it faces significant challenges in scalability and consistency due to reliance on human judgment. Multi-turn dialogue frameworks, despite better capturing interactive complexity, lack specific mechanisms for eliciting and evaluating teaching effectiveness. These limitations particularly impact teaching evaluation: methodologies neither capture teachers’ active role in questioning, assessment, and real-time adaptation, nor provide scalable solutions for evaluating teaching quality.
To address these challenges, we propose EducationQ, a novel multi-agent dialogue framework that incorporates formative assessment into the evaluation of LLMs’ teaching capabilities. Formative assessment—a continuous process of assessing learner progress, identifying gaps, and adjusting teaching strategies William (2011)—is essential for personalized instruction Miao et al. (2021). It bridges the gap between students’ current abilities and potential, enhances learning outcomes, and promotes educational equity through AI Pardo et al. (2019); Ruiz-Primo and Furtak (2007); U.S. Department of Education, Office of Educational Technology (2023); Allal and Pelgrims Ducrey (2000). In real-world classrooms, these principles are commonly manifested as informal formative assessments (IFAs) during instructional dialogues, where teachers pose questions, assess student understanding, and provide timely feedback and guidance Sezen-Barrie and Kelly (2017); Guskey (2005).
The EducationQ framework models these interactions through a triadic system of teacher, student, and evaluator agents, simulating cyclical teacher-student interaction. This design captures teachers’ agency in employing diverse strategies and navigating complex educational contexts while enabling automated evaluation of dialogue quality. To support this framework, we curated a robust dataset of 1,498 questions from established benchmarks GPQA and MMLU-Pro, spanning diverse disciplines and difficulty levels. We employed a mixed-methods approach to comprehensively evaluate LLMs’ teaching capabilities. This approach quantifies teaching effectiveness from an outcome-aligned perspective Gitomer and Duschl (2007) by measuring student learning gains, while analyzing pedagogical strategies using an automated qualitative evaluator agent with strong alignment (78% agreement) with human education experts.
@@
  • A multi-agent dialogue methodology for simulating and analyzing teaching interactions.
@@ 2.1 LLM Evaluation
Task-oriented performance benchmarks like MMLU Hendrycks et al. (2021a), MMLU-Pro Wang et al. (2024b), and GPQA Rein et al. (2023) employ closed-ended questions to evaluate domain knowledge and reasoning abilities. MATH Hendrycks et al. (2021b) examines mathematical reasoning, while HumanEval Chen et al. (2021) tests programming capabilities. MathChatLiang et al. (2024) evaluates LLMs on multi-turn mathematical interactions, revealing that models struggle with sustained reasoning across multiple dialogue turns.
@@ 2.3 LLM-Based Student Modeling
Recent advancements have demonstrated the potential of LLMs to simulate nuanced human cognitive processes and behaviors Park et al. (2023), including emulating reasoning pathways Wei et al. (2022); Kojima et al. (2022), and mimicking human-like reflective thinking Weng et al. (2023). This burgeoning ability to model human-like thought and action has naturally led to applying LLMs to simulate student behavior and interactions within educational contexts. Xu & Zhang (2023) investigated the feasibility of using generative students to test educational materials, while Markel et al. (2023) employed LLMs to simulate student dialogues for teacher training. Further, Lu & Wang (2024) found that profile-based generative students can closely mirror human student performance in MCQ responses. Jin et al. (2025) proposed TeachTune, a framework generating pedagogical agent dialogues with diverse simulated student profiles for human evaluation, complementing our automated fixed-student-model assessment approach.
@@ 3 Dataset
We constructed our evaluation datasets, as summarized in Table 1, by systematically curating questions from two well-established benchmarks: GPQA (n=448), featuring domain expert-authored questions, and MMLU-Pro (n=12,032), containing reasoning-intensive questions with enhanced robustness through 10-option design across 14 categories. These datasets span undergraduate to PhD-level content, providing a rigorous foundation for evaluating teaching capabilities. The resulting dataset benefits from the well-established credibility of source benchmarks, proven stability of their evaluation protocols, and comprehensive representation across multiple teaching dimensions.
@@
Figure 2: Dataset distribution across 13 academic disciplines and 10 difficulty levels. Left: original MMLU-Pro shows uneven distribution across subjects and difficulty levels; Right: MMLU-Pro Stratified with balanced sampling approach ensuring equal representation.
@@ 4.2 Teacher Agent
To prevent direct answer disclosure, we implement strict data flow controls that technically prevent teacher agents from accessing any of the question options. This enforced constraint, combined with explicit prompt instructions, ensures teachers must guide learning without revealing answers.
@@ 4.3 Evaluator Agent
The evaluator agent is prompted (see Appendix A.3) as an education assessment expert well-versed in pedagogical theory and practice. It validates the dialogues and evaluates teaching according to specified dimensions and compares teacher performances to determine superior approaches.
The qualitative analysis framework comprises 17 distinct scoring dimensions, including teacher-focused metrics (questioning, assessment, feedback) and student-impact measures (metacognitive reflection, knowledge dimension, etc.) Krathwohl (2002); Looney (2011); Wilen (1987); Wass and and (2014). Given the exploratory nature of this component, we did not address the dimensional overlap.
@@ 4.5 Pre-Test
The pre-test establishes the student agent’s initial knowledge baseline while providing teachers with preliminary insights through chain-of-thought reasoning patterns. To ensure broader applicability and stability, we conducted pre-test evaluation following the official MMLU-Pro and GPQA Diamond benchmark protocols and parameters.
@@ 4.8 Content Boundary Design
As mentioned in the Teacher Agent section, to prevent direct answer disclosure, we implemented the following technical constraints: (a) Teacher agents cannot access any of the answer options, relying solely on student reasoning patterns and correctness judgments for guidance. (b) Students cannot access pre-test correctness judgments during dialogue, learning exclusively through teacher interaction. (c) Students retain access to complete question content including options, enabling learning through experience association.
@@ 4.9 Interaction Parameters and Constraints
Ablation studies identified optimal parameters of 150 tokens per turn across five rounds, balancing effectiveness and efficiency. Increasing token limits to 250 yielded no significant learning gains, while reducing teacher dialogue to 70-100 tokens degraded teaching performance. Doubling rounds to 10 (with halved student token limits) increased computational costs without surpassing the effectiveness of the 5-round, 150-token configuration. In final experiments, teacher responses averaged 73.6 tokens, with student responses averaging 260 tokens.
@@ 7.2 Framework Robustness and Stability
Table 3: Stability Study of the Multi-agent Framework.
@@
Table 4: Teaching Performance Comparison of Large Language Models.
@@ 7.5 Teaching Stability Analysis
Through analysis of the Positive-Negative Impact Ratio (PNIR, Equation 2, lower is better), we identified significant variations in teaching stability across models. Llama 3.1 70B Instruct demonstrated exceptional stability, generating only 36 negative cases against 200 positive improvements (PNIR = 0.18). While Gemini 1.5 Pro 002 achieved comparable positive cases (188), its higher PNIR of 0.40 indicated greater performance volatility. OpenAI o1-mini and Llama 3.1 405B Instruct maintained moderate stability (PNIR = 0.25 and 0.28 respectively). These findings suggest that high teaching effectiveness and high stability can coexist.
@@ 8.1 Teaching Strategy Analysis: Expert-Annotated Case Studies
Human education experts confirmed that learning gains resulted from diverse and sophisticated teaching strategies rather than direct answer disclosure. The experts identified and annotated clear evidence of pedagogical techniques. Our expert-annotated case studies provide detailed evidence of how different LLMs implement distinct teaching strategies and demonstrate varying pedagogical skills and explain why models with superior general capabilities might underperform in educational interactions.
@@ 9 Conclusion
These findings suggest a fudamental rethinking of educational LLM development: prioritizing specialized teaching capabilities over general model scaling. The significant performance variations also highlight the inadequacy of traditional LLM metrics for predicting teaching effectiveness, underscoring the critical need for specialized, interaction-based evaluation frameworks like EducationQ to guide the development of effective AI in education.
@@
While our primary metric, ALG, objectively measures learning outcomes, the automated qualitative analysis, though validated with 78% human agreement, captures a predefined set of pedagogical dimensions. Further research could explore even broader qualitative aspects or integrate them more directly into a composite teaching score, if deemed necessary.
@@ A.1 Student Agent Configuration
The student agent uses a consistent prompting template designed to simulate authentic student learning behavior, including expressing uncertainty and focusing on problem analysis. Below we detail the exact configuration and prompting templates used in our experiments for the student agent.
@@ A.1.3 Response Format
For dialogue interactions, The student agent’s interaction involves receiving context from previous rounds and generating a new response:
@@
Note: [SYSTEM_MESSAGE] refers to the template shown in Appendix A.1.2. [PRE_TEST_INFO] is included in the context if the configuration parameter (include_pretest_info) is enabled.
@@ A.2.3 Pre-test Information Format
Note: The [CORRECTNESS] is either "Correct." or "Incorrect." based on the student’s selection.
@@ A.3.3 Evaluation Dimensions
The evaluator assesses effectiveness based on dimensions specific to each analytical perspective:
@@ A.3.4 Evaluation Input Format (Example: Holistic Interaction Analysis)
Note: For Teacher/Student-centric analysis, only the questions/answers are included.
@@ A.3.5 Evaluation Output Format (Each Dimension)
For each dimension within an analysis task, the evaluator provides a JSON object structured as follows (refer to Appendix B.3 for scoring guidelines):
@@ A.3.6 Comparative Analysis Output Format
Note: The code includes logic to de-anonymize the teacher_a/teacher_b keys and the final choice based on the actual models being compared.
The implementation details, can be found in the EvaluatorLLM class.
@@ B.2 Quality Control and Response Validation
Mechanisms ensure the quality and validity of LLM responses during generation and processing:
@@ B.3 Evaluator Scoring Guidelines
Scoring performed by the Evaluator agent (Appendix A.3.2) follows a standardized 1-10 scale:
@@ Appendix C Model Specifications
  • Note: "-" indicates unspecified information; Context window sizes are in tokens;
  • Org.: Organization (model developer); Provider: serving platform.
@@ 14 GPQA
Table 8: Dataset Distribution by Source and Discipline
@@ Appendix E Teaching Performance Ranking by Subject
Tables 9 through 21 present the detailed Absolute Learning Gain ($\Delta$, in %) for each teacher model within 13 academic disciplines.